# RepairLLaMA: Efficient Representations and Fine-Tuned Adapters for Program Repair

### v5

@@ 1 Introduction
Fine-tuning has the potential to learn domain-specific representations and leverage signals in order to generate high-quality patches. The drawback is that it requires substantial computational resources and high-quality training data. Prompt engineering and agent-based approaches mostly consist of ad hoc human-developed prompts and workflows to instruct models for specific tasks and are bounded by the limited size of context windows. Both lines make different trade-offs in applying LLMs to automated bug fixing. In this paper, we make an important contribution in the space of fine-tuning LLMs for program repair.
@@
  • We demonstrate that parameter-efficient fine-tuning performs better than full-parameter fine-tuning in the context of program repair. The “repair adapters” of RepairLLaMA are training-efficient, and achieve state-of-the-art repair performance across three benchmarks, Defects4J, HumanEval-Java, and GitBug-Java, outperforming even GPT-4.
@@
2 RepairLLaMA: Efficient Fine-Tuning for Program Repair
2.1 Overview
@@
In the first stage of RepairLLaMA, the core choices are made, namely: 1) the initial pre-trained model (subsection 2.3); 2) the input code representation and output code representation (subsection 2.4); and 3) the fine-tuning dataset (subsection 2.5). These choices are all important and are further discussed in the remainder of this section.
@@ 2.3 Choice of the Initial LLM
In subsection 3.2 we instantiate those criteria in the context of functional program repair for Java.
2.4 Choice of Code Representations
@@
2.4.1 Representation of Fault Localization
@@
2.4.2 Input Representation Space
@@
2.4.3 Output Representation Space
@@
Output representations in APR correspond to the representation of the synthesized fixed code. A natural output representation is a diff over the buggy code, aka a patch. As discussed in subsubsection 2.4.2, fine-tuning is required to adapt an LLM to generate such task-specific outputs. To study the output representation space, we design four output representations tailored to APR (Figure 3):
@@
2.4.4 Input/Output Representation Pairs
@@
2.5 Choice of Fine-Tuning Dataset
@@
2.6 Program Repair Adapters for LLMs
@@ 2.7 Inference Time
3 Experimental Methodology
3.1 Research Questions
@@
3.2 Implementation
Model to Fine-Tune Per the criteria of subsection 2.3, we choose CodeLlama-7b [24] as our initial LLM. CodeLLaMA is a publicly available LLM released in 2023 and is trained on 500B code tokens. Per the experiments reported in [24], CodeLLaMA outperforms GPT-3.5 on two code generation benchmarks.
@@ 3.2 Implementation
Evaluation Benchmark We select three Java benchmarks for our evaluation: Defects4J [22], HumanEval-Java [8], and GitBug-Java [23]. Following recent related work [34, 11, 14], we scope our evaluation to single-function bugs, as defined in subsection 2.2. Defects4J comprises 835 real-world bugs from 17 open-source Java projects, from which we identify 488 single-function bugs. HumanEval-Java is a bug benchmark containing artificial bugs inserted in HumanEval [35] Java programs. HumanEval-Java contains 162 single-function bugs. GitBug-Java is a bug benchmark of recent bugs, collected from the 2023 commit history of 55 open-source repositories, comprising 90 single-function bugs. Contrary to Defects4J, GitBug-Java suffers from less data leakage in the pre-training data since all bugs are much more recent than Defects4J. Notably, GitBug-Java exclusively contains bugs from after the training data cutoff date of all models used in our experiments: CodeLlama-7b (September 2022) [24], gpt-4-0613 (September 2021)  [^2] https://platform.openai.com/docs/models, and gpt-3.5-turbo-0613 (September 2021) \@footnotemark.
@@ 3.3 Patch Assessment
3.4 Methodology for RQ1
The objective of RQ1 is to investigate the most effective code representations for fine-tuning an LLM for program repair. While existing research has delved into the utility of LLMs for program repair, the impact of the code representations, such as their realism, has been overlooked. It is known that variations in code representations may yield substantial differences in performance for fine-tuned LLMs [25]. Consequently, in RQ1, we empirically evaluate 6 realistic code representation pairs presented in 2.4 and measure their performance.
We fine-tune an LLM as described in 3.2. We prompt the model to generate 10 patches for each bug using beam search decoding. We then evaluate the generated patches as outlined in subsection 3.3, to measure the effectiveness of each code representation. We prompt the non fine-tuned CodeLLaMA-7B as a baseline.
@@ 3.5 Methodology for RQ2
Baseline. We consider four baselines in RQ2. One is the baseline considered in RQ1. The second one comprises the full-parameter fine-tuned version of RepairLLaMA, a powerful but naive approach to fine-tuning a large language model for program repair. Here, we use the same hyper-parameters as in LoRA fine-tuning, apart from a lower learning rate of 2e-5. We employ the same statistical significance test as described in subsection 3.4 when comparing RepairLlama against these two baselines.
The third is Jiang et al.’s work [8]. We compare against the best performing model reported, the fine-tuned version of Incoder-6B [38]. The last one is RAP-Gen [10], a state-of-the-art fine-tuned LLM for program repair, which uses fix patterns retrieved from a codebase of previous bug-fix pairs to augment the buggy input.
3.6 Methodology for RQ3
@@
First, we zero-shot prompt gpt-3.5-turbo-0613 and gpt-4-0613 to generate 10 patches for each bug. The prompt is shown in Figure 4, it is built by integrating the best code representation with an effective zero-shot prompt from Zhang et al. [39]. The prompt instructs the LLM to generate the fixed code chunk to replace the <FILL_ME> token. Note that the same information signals are provided to both RepairLLaMA models and the GPT models, allowing for a fair comparison of the repair capabilities of the different models. This includes providing the same source code context, as well as the same fault localization information as input. We employ the same statistical significance test as described in subsection 3.4. We utilize OpenAI’s official APIs to call gpt-3.5-turbo-0613 and gpt-4-0613 to conduct our experiments on Dec. 1, 2023, except for the GitBug-Java experiments which are conducted on gpt-3.5-turbo-0125 on Nov. 17, 2024.
4 Experimental Results
4.1 Results of RQ1 (Code Representations for Fine-Tuning)
@@
{tblr}
row2 = c, column2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4c, cell17 = c=4c, cell111 = c=4c, cell33 = r, cell34 = r, cell35 = r, cell36 = r, cell37 = r, cell38 = r, cell39 = r, cell310 = r, cell311 = r, cell312 = r, cell313 = r, cell314 = r, cell43 = r, cell44 = r, cell45 = r, cell46 = r, cell47 = r, cell48 = r, cell49 = r, cell410 = r, cell411 = r, cell412 = r, cell413 = r, cell414 = r, cell53 = r, cell54 = r, cell55 = r, cell56 = r, cell57 = r, cell58 = r, cell59 = r, cell510 = r, cell511 = r, cell512 = r, cell513 = r, cell514 = r, cell63 = r, cell64 = r, cell65 = r, cell66 = r, cell67 = r, cell68 = r, cell69 = r, cell610 = r, cell611 = r, cell612 = r, cell613 = r, cell614 = r, cell73 = r, cell74 = r, cell75 = r, cell76 = r, cell77 = r, cell78 = r, cell79 = r, cell710 = r, cell711 = r, cell712 = r, cell713 = r, cell714 = r, cell83 = r, cell84 = r, cell85 = r, cell86 = r, cell87 = r, cell88 = r, cell89 = r, cell810 = r, cell811 = r, cell812 = r, cell813 = r, cell814 = r, cell93 = r, cell94 = r, cell95 = r, cell96 = r, cell97 = r, cell98 = r, cell99 = r, cell910 = r, cell911 = r, cell912 = r, cell913 = r, cell914 = r, cell103 = r, cell104 = r, cell105 = r, cell106 = r, cell107 = r, cell108 = r, cell109 = r, cell1010 = r, cell1011 = r, cell1012 = r, cell1013 = r, cell1014 = r, vline2-15 = 1, vline4-15 = 2, vline2-15 = 2-10, hline1,3-11 = 2-14, hline2 = 3-14, & Code Representations Defects4J v2
(488 bugs)
HumanEval-Java
(162 bugs)
GitBug-Java
@@ 8 IR4 x OR2 (RepairLLaMA) 195 124 125 144 118 75 82 109 25 15 16 20
In RQ1, we investigate the most effective code representations for fine-tuning an LLM for program repair. The results of the evaluation are presented in Table II, which shows the effectiveness of each code representation setting on both test benchmarks. The table is structured as follows: the first column displays the code representations, the second to fourth meta-columns show the repair effectiveness results on, respectively, the single-function subsets of Defects4J v2, HumanEval-Java, and GitBug-Java. Recall that the repair effectiveness evaluation is measured by the four patch assessment metrics described in subsection 3.3.
@@
4.2 Results of RQ2 (Parameter-Efficient Fine-Tuning vs. Full Fine-Tuning)
Table III: Repair effectiveness of RepairLLaMA compared with fully fine-tuned models (RQ2). RepairLLaMA, trained with parameter-efficient fine-tuning, outperforms all comparative models on Defects4J, HumanEval-Java, and GitBug-Java.
{tblr}
cells = r, row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell31 = c, cell41 = c, cell51 = c, vlines, hline1,3-6 = -, hline2-3 = 2-13, Model & Defects4J v2
(488 bugs)
HumanEval-Java
(162 bugs)
GitBug-Java
@@
IR3 x OR2 (no fine-tuning) 131 52 70 83 107 71 81 103 17 8 9 12
IR4 x OR2 (full fine-tuning) 146 66 84 98 109 74 83 100 21 11 11 13
IR4 x OR2 (RepairLLaMA) 195 124 125 144 118 75 82 109 25 15 16 20
In RQ2, we study how parameter-efficient fine-tuning compares against basic full-parameter fine-tuning. Recall most of the closely related work in APR [8, 10, 41, 9, 12, 5] use full-parameter fine-tuning. On the contrary, the key novelty of RepairLLaMA is that all models are fine-tuned using LoRA, a parameter-efficient fine-tuning technique that optimizes only a small adapter (approx. 4M parameters) instead of the whole LLM (approx. 7B parameters, a reduction of approx. 1600x). This allows the model to 1) be fine-tuned with less GPU memory, and 2) potentially reduce overfitting. In RQ2, we compare RepairLLaMA, built with the best code representation in RQ1, with its full-parameter fine-tuning version.
Table III presents the results of RQ2. The table reads as follows. The first column presents the model. The second to fourth meta-columns show the repair effectiveness results on, respectively, Defects4J v2, HumanEval-Java, and GitBug-Java.
@@
The gain in performance is clear, and it is actually a double win because RepairLLaMA requires fewer resources. Possibly, the fully fine-tuned models suffer from overfitting due to the limited fine-tuning data, while parameter-efficient fine-tuning helps prevent overfitting since it only requires optimizing a small part of network weights. The model size constraints of LoRA appear to act as implicit regularizers in this experiment. These results confirm the findings of Fu et al. [21]: parameter-efficient fine-tuning helps mitigate overfitting by leveraging trainable parameter sparsity and reducing the generalization error bound. Yet, a balance is needed to prevent underfitting due to insufficient trainable parameters.
When compared with the best fully fine-tuned model from Jiang et al. [8], which keeps the LLMs’ original pre-training format but includes the buggy lines as comments, RepairLLaMA correctly repairs 75 (144 vs. 69) Defects4J bugs more and 48 (118 vs. 70) HumanEval-Java bugs more. Notably, RepairLLaMA achieves these results without access to perfect bug localization information, which is assumed and incorporated in [8].
@@
Answer to RQ2: RepairLLaMA’s parameter-efficient fine-tuning outperforms full fine-tuning, with external validity over three datasets (Defects4J, HumanEval-Java, and GitBug-Java). The fair experimental comparison is a clear-cut higher performance: 44 additional bugs are correctly repaired in Defects4J thanks to parameter-efficient fine-tuning. Overall, our paper is the first to demonstrate the advantage of parameter-efficient fine-tuning in the context of automated program repair, achieving strong state-of-the-art results. Beyond program repair, we note that parameter-efficient fine-tuning is feasible within academic labs’ budget and hardware, while still working with powerful multi-million dollar trained large models.
4.3 Results of RQ3 (RepairLLaMA vs. ChatGPT-based APR)
@@
{tblr}
cells = r, row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell31 = c, cell41 = c, cell51 = c, vlines, hline1,4-6 = -, hline2-3 = 1-13, Model & Defects4J v2
(488 bugs)
HumanEval-Java
(162 bugs)
GitBug-Java
@@
Table IV shows the repair effectiveness of RepairLLaMA compared with ChatGPT-based APR techniques. The first column indicates the model name. The second, and third show the results on two different benchmarks. Each benchmark is evaluated following the four patch assessment metrics described in subsection 3.3.
@@
5 Discussion
5.1 Sampling Candidate Patches

### v6

@@ I Introduction
Fine-tuning has the potential to learn domain-specific representations and leverage signals in order to generate high-quality patches. The drawback is that it requires substantial computational resources and high-quality training data. Prompt engineering and agent-based approaches mostly consist of ad hoc human-developed prompts and workflows to instruct models for specific tasks and are bounded by the limited size of context windows. While both approaches offer different trade-offs, we argue that fine-tuning is crucial for high-quality automated program repair because: (1) it allows the model to internalize repair patterns from a supervised corpus of bug fixes, (2) it enables learning of task-specific repair representations rather than relying on manual prompt engineering steps, and (3) it provides opportunities to incorporate repair-specific information during training. Our paper is a key contribution in the space of fine-tuning LLMs for program repair.
@@
  • We demonstrate that parameter-efficient fine-tuning performs competitively with full-parameter fine-tuning in the context of program repair. The “repair adapters” of RepairLLaMA are training-efficient, and achieve state-of-the-art repair performance across three benchmarks, Defects4J, HumanEval-Java, and GitBug-Java, outperforming even GPT-4.
@@
II RepairLLaMA: Efficient Fine-Tuning for Program Repair
II-A Overview
@@
In the first stage of RepairLLaMA, the core choices are made, namely: 1) the initial pre-trained model (subsection II-C); 2) the input code representation and output code representation (subsection II-D); and 3) the fine-tuning dataset (subsection II-E). These choices are all important and are further discussed in the remainder of this section.
@@ II-C Choice of the Initial LLM
In subsection III-B we instantiate those criteria in the context of functional program repair for Java.
II-D Choice of Code Representations
@@
II-D1 Representation of Fault Localization
@@
II-D2 Input Representation Space
@@
II-D3 Output Representation Space
@@
Output representations in APR correspond to the representation of the synthesized fixed code. A natural output representation is a diff over the buggy code, aka a patch. As discussed in subsubsection II-D2, fine-tuning is required to adapt an LLM to generate such task-specific outputs. To study the output representation space, we design four output representations tailored to APR (Figure 3):
@@
II-D4 Input/Output Representation Pairs
@@
II-E Choice of Fine-Tuning Dataset
@@
II-F Program Repair Adapters for LLMs
@@ II-G Inference Time
III Experimental Methodology
III-A Research Questions
@@
III-B Implementation
Model to Fine-Tune Per the criteria of subsection II-C, we choose CodeLlama-7b [24] as our initial LLM. CodeLLaMA is a publicly available LLM released in 2023 and is trained on 500B code tokens. Per the experiments reported in [24], CodeLLaMA outperforms GPT-3.5 on two code generation benchmarks.
@@ III-B Implementation
Evaluation Benchmark We select three Java benchmarks for our evaluation: Defects4J [22], HumanEval-Java [8], and GitBug-Java [23]. Following recent related work [34, 11, 14], we scope our evaluation to single-function bugs, as defined in subsection II-B. Defects4J comprises 835 real-world bugs from 17 open-source Java projects, from which we identify 488 single-function bugs. HumanEval-Java is a bug benchmark containing artificial bugs inserted in HumanEval [35] Java programs. HumanEval-Java contains 162 single-function bugs. GitBug-Java is a bug benchmark of recent bugs, collected from the 2023 commit history of 55 open-source repositories, comprising 90 single-function bugs. Contrary to Defects4J, GitBug-Java suffers from less data leakage in the pre-training data since all bugs are much more recent than Defects4J. Notably, GitBug-Java exclusively contains bugs from after the training data cutoff date of all models used in our experiments: CodeLlama-7b (September 2022) [24], gpt-4-0613 (September 2021)  [^2] https://platform.openai.com/docs/models, and gpt-3.5-turbo-0613 (September 2021) \@footnotemark.
@@ III-C Patch Assessment
III-D Methodology for RQ1
The objective of RQ1 is to investigate the most effective code representations for fine-tuning an LLM for program repair. While existing research has delved into the utility of LLMs for program repair, the impact of the code representations, such as their realism, has been overlooked. It is known that variations in code representations may yield substantial differences in performance for fine-tuned LLMs [25]. Consequently, in RQ1, we empirically evaluate 6 realistic code representation pairs presented in II-D and measure their performance.
We fine-tune an LLM as described in III-B. We prompt the model to generate 10 patches for each bug using beam search decoding. We then evaluate the generated patches as outlined in subsection III-C, to measure the effectiveness of each code representation. We prompt the non fine-tuned CodeLLaMA-7B as a baseline.
@@ III-E Methodology for RQ2
Baseline. We consider two baseline base models in RQ2: CodeLLaMA-7B and deepseek-coder-6.7b-base [38]. To study the difference between parameter-efficient fine-tuning and full-parameter fine-tuning, we fine-tune both models with both approaches. Here, we use the same hyper-parameters as in LoRA fine-tuning described in RQ1. We use a learning rate of 2e-5 for full-parameter-fine-tuning. We employ the same statistical significance test as described in subsection III-D when comparing RepairLlama against these baselines.
Additionally, we benchmark our approach against the highest-performing model documented by Jiang et al. [34]—specifically, the fine-tuned variant of Incoder-6B [39]. We also compare with RAP-Gen [10], a state-of-the-art model for automated program repair that enhances processing of buggy inputs by leveraging fix patterns retrieved from an extensive repository of historical bug-fix pairs.
III-F Methodology for RQ3
@@
First, we zero-shot prompt gpt-3.5-turbo-0613 and gpt-4-0613 to generate 10 patches for each bug. The prompt is shown in Figure 4, it is built by integrating the best code representation with an effective zero-shot prompt from Zhang et al. [40]. The prompt instructs the LLM to generate the fixed code chunk to replace the <FILL_ME> token. Note that the same information signals are provided to both RepairLLaMA models and the GPT models, allowing for a fair comparison of the repair capabilities of the different models. This includes providing the same source code context, as well as the same fault localization information as input. We employ the same statistical significance test as described in subsection III-D. We utilize OpenAI’s official APIs to call gpt-3.5-turbo-0613 and gpt-4-0613 to conduct our experiments on Dec. 1, 2023, except for the GitBug-Java experiments which are conducted on gpt-3.5-turbo-0125 on Nov. 17, 2024.
IV Experimental Results
IV-A Results of RQ1 (Code Representations for Fine-Tuning)
@@
Code Representations Defects4J v2
(488 bugs) HumanEval-Java
(162 bugs) GitBug-Java
@@ 8 IR4 x OR2 (RepairLLaMA) 195 124 125 144 118 75 82 109 25 15 16 20
In RQ1, we investigate the most effective code representations for fine-tuning an LLM for program repair. The results of the evaluation are presented in subsection IV-A, which shows the effectiveness of each code representation setting on both test benchmarks. The table is structured as follows: the first column displays the code representations, the second to fourth meta-columns show the repair effectiveness results on, respectively, the single-function subsets of Defects4J v2, HumanEval-Java, and GitBug-Java. Recall that the repair effectiveness evaluation is measured by the four patch assessment metrics described in subsection III-C.
@@
IV-B Results of RQ2 (Parameter-Efficient Fine-Tuning vs. Full Fine-Tuning)
Table III: Repair effectiveness of parameter-efficient fine-tuned models compared with fully fine-tuned models (RQ2). RepairLLaMA, trained with parameter-efficient fine-tuning, outperforms its variants on Defects4J, HumanEval-Java, and GitBug-Java.
Base
Model Fine-Tuning
Method Defects4J v2
(488 bugs) HumanEval-Java
(162 bugs) GitBug-Java
@@
codellama-7b - 131 $52$ 70 83 107 71 81 103 17 8 9 12
codellama-7b FFT 146 $66$ 84 98 109 74 83 100 21 11 11 13
codellama-7b LoRA 195 124 125 144 118 75 82 109 25 15 16 20
deepseek-coder-6.7b Base 147 $84$ 88 104 113 68 83 110 21 11 11 13
deepseek-coder-6.7b FFT 185 116 123 138 129 96 101 119 26 15 15 19
deepseek-coder-6.7b LoRA 181 $110$ 113 128 134 99 107 124 19 12 12 13
In RQ2, we study how parameter-efficient fine-tuning compares against basic full-parameter fine-tuning. Recall most of the closely related work in APR [8, 10, 42, 9, 12, 5] use full-parameter fine-tuning. On the contrary, the key novelty of RepairLLaMA is that all models are fine-tuned using LoRA, a parameter-efficient fine-tuning technique that optimizes only a small adapter (approx. 4M parameters) instead of the whole LLM (approx. 7B parameters, a reduction of approx. 1600x). This allows the model to 1) be fine-tuned with less GPU memory, and 2) potentially reduce overfitting. In RQ2, we compare RepairLLaMA, built with the best code representation in RQ1, with its full-parameter fine-tuning version. Additionally, we also study the same fine-tuning methods with a different base model, deepseek-coder-6.7b, to increase external validity.
Table III presents the results of RQ2. The table reads as follows. The first column presents the base model. The second column presents the fine-tuning model. The third to fifth meta-columns show the repair effectiveness results on, respectively, Defects4J v2, HumanEval-Java, and GitBug-Java.
@@
The gain in performance is clear, and it is actually a double win because RepairLLaMA requires fewer resources. Possibly, the fully fine-tuned models suffer from overfitting due to the limited fine-tuning data, while parameter-efficient fine-tuning helps prevent overfitting since it only requires optimizing a small part of the network weights. The model size constraints of LoRA appear to act as implicit regularizers in this experiment. These results confirm the findings of Fu et al. [21]: parameter-efficient fine-tuning helps mitigate overfitting by leveraging trainable parameter sparsity and reducing the generalization error bound. Yet, a balance is needed to prevent underfitting due to insufficient trainable parameters.
This phenomenon is not observed for deepseek-coder-6.7b. For this model, fine-tuning clearly improves over the base model. The best fine-tuned model, for each benchmark, repairs 34 (138 vs. 104) more Defects4J, 24 (107 vs. 83) more HumanEval-Java, and 6 (19 vs. 13) more GitBug-Java bugs than the base model. The full fine-tuned version outperforms LoRA in Defects4J and GitBug-Java, while the LoRA version outperforms in HumanEval-Java. There is a statistically significant difference between both the fully fine-tuned ($p\approx 2.72e-05$) and LoRA versions ($p\approx 4.21e-04$) of deepseek-coder-6.7b when compared with the baseline model, meaning that both methods perform a statistically significant improvement over the baseline model. These results suggest that the effect of parameter-efficient fine-tuning varies across models. It can under- or over-perform depending on the model and benchmark, always with the advantage of smaller resource requirements.
When compared with the best fully fine-tuned model from Jiang et al. [8], which keeps the LLMs’ original pre-training format but includes the buggy lines as comments, RepairLLaMA correctly repairs 25 (95 vs. 69) Defects4J bugs more, when applied to the same dataset considered in Jiang et al. and 39 (109 vs. 70) HumanEval-Java bugs more. In addition, RepairLLaMA correctly repairs 49 more Defects4J multi-hunk bugs. Notably, RepairLLaMA achieves these results without access to perfect bug localization information, which is assumed and incorporated in [8].
@@
Figure 5 shows a multi-location bug from Math-86 (Defects4J v2) where RepairLLaMA successfully generates an exact match patch by fixing two distant buggy locations simultaneously. This example clearly demonstrates the limitations of both comparison approaches: Jiang et al.’s approach can only handle single-chunk bugs, and its representation is inadequate for addressing bugs that span multiple locations, as shown in this example. RAG-Gen, while theoretically capable, requires perfect fault localization information, which is impractical in real-world scenarios. In contrast, RepairLLaMA overcomes these limitations by tailoring effective code representation that can handle complex multi-location bugs without requiring perfect fault localization information.
Answer to RQ2: The efficacy of fine-tuning is clear, yet depends on the base model. RepairLLaMA’s parameter-efficient fine-tuning outperforms full fine-tuning of codellama-7b, with external validity over three benchmarks, while results for deepseek-coder-6.7b show similar improvements regardless of the fine-tuning method. Overall, our paper is the first to demonstrate the benefits of parameter-efficient fine-tuning in the context of automated program repair, achieving strong state-of-the-art results. Beyond program repair, we note that parameter-efficient fine-tuning is feasible within academic labs’ budget and hardware, while still working with powerful multi-million dollar trained large models.
IV-C Results of RQ3 (RepairLLaMA vs. ChatGPT-based APR)
@@
Model Defects4J v2
(488 bugs) HumanEval-Java
(162 bugs) GitBug-Java
@@
Table IV shows the repair effectiveness of RepairLLaMA compared with ChatGPT-based APR techniques. The first column indicates the model name. The second, and third show the results on two different benchmarks. Each benchmark is evaluated following the four patch assessment metrics described in subsection III-C.
@@
V Discussion
V-A Sampling Candidate Patches